OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step. I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2

OpenStudy (anonymous):

2x + 4 = 3x – 11 3x + 3 = 51 2x + 1/ 3 = x + 6 / 2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (rsadhvika):

Start by solving first equation

OpenStudy (anonymous):

2x + 4 = 3x - 11

OpenStudy (anonymous):

add 11 to both sides?

OpenStudy (anonymous):

wait no, subtract 4 first i think

OpenStudy (rsadhvika):

``` I. 2x + 4 = 3x – 11 ``` yeah u can subtract 4 both sides. wat do u get

OpenStudy (anonymous):

2x + 4 = 3x - 11 -4 -4 2x = 3x - 7

OpenStudy (rsadhvika):

next, subtract 3x both sides

OpenStudy (rsadhvika):

wait a sec

OpenStudy (rsadhvika):

2x + 4 = 3x - 11 -4 -4 2x = 3x - 15

OpenStudy (anonymous):

2x + 4 = 3x - 11 -4 -4 2x = 3x - 7 -3x -3x -1x = 7

OpenStudy (rsadhvika):

u shud get that, -11-4 = -15

OpenStudy (anonymous):

alright -1x = 15

OpenStudy (rsadhvika):

2x + 4 = 3x - 11 -4 -4 2x = 3x - 15 -3x -3x -x = -15 x = 15

OpenStudy (rsadhvika):

thats the solution for first equation

OpenStudy (anonymous):

-1x = -15 /-1 /-1 x = 15?

OpenStudy (rsadhvika):

yup, and u need to explain each step. just add lines after each step, saying things like :- subtract 4 both sides... etc

OpenStudy (anonymous):

alright gimme 1 sec to write this down :)

OpenStudy (anonymous):

alright number 2, 3x + 3 = 51

OpenStudy (anonymous):

I figured out 2, just need help on 3

OpenStudy (anonymous):

2 is x = 16?

OpenStudy (rsadhvika):

for equation 2 :- ``` 3x + 3 = 51 ``` subtract 3 both sides 3x = 48 divide 3 both sides x = 16

OpenStudy (anonymous):

yep thats what I did :D

OpenStudy (rsadhvika):

for equation 3 :- ``` III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2 ``` \(\large \frac{2x+1}{3} = \frac{x+6}{2}\) cross multiply 2(2x+1) = 3(x+6) distribute 4x + 2 = 3x +18 subtract 3x both sides x + 2 = 18 subtract 2 both sides x = 16